### v3

# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

Abstract
Accurate forecasting of epidemic infection trajectories is crucial for safeguarding public health. However, limited data availability during emerging outbreaks and the complex interaction between environmental factors and disease dynamics present significant challenges for effective forecasting. In response, we introduce CAPE, a novel epidemic pre-training framework designed to harness extensive disease datasets from diverse regions and integrate environmental factors directly into the modeling process for more informed decision-making on downstream diseases. Based on a covariate adjustment framework, CAPE utilizes pre-training combined with hierarchical environment contrasting to identify universal patterns across diseases while estimating latent environmental influences. We have compiled a diverse collection of epidemic time series datasets and validated the effectiveness of CAPE under various evaluation scenarios, including full-shot, few-shot, zero-shot, cross-location, and cross-disease settings, where it outperforms the leading baseline by an average of 9.9% in full-shot and 14.3% in zero-shot settings. The code will be released upon acceptance.
Machine Learning, ICML
1 Introduction
Infectious disease outbreaks consistently challenge public health systems, affecting both individual well-being and economic stability (Nicola et al., 2020). Effective management of these outbreaks hinges on accurate epidemic forecasting, which involves predicting future incidences like infection cases and hospitalizations (Liu et al., 2024b; Wan et al., 2024; Adhikari et al., 2019). Over the years, various models have been developed to address this need. These include mechanistic models like SIR (Cooper et al., 2020) and statistical models like ARIMA (Sahai et al., 2020; Kontopoulou et al., 2023), as well as advanced machine learning methods such as LSTM and GRU (Shahid et al., 2020), which have proven instrumental in forecasting disease spread and supporting informed public health decision-making.
Despite the advancements, current models are typically trained for specific diseases within particular geographic regions, limiting their ability to integrate insights from diverse sources spanning multiple pathogens and spatial contexts. This narrow focus can impede a comprehensive understanding of disease dynamics and the design of effective outbreak responses, especially during novel or emergent outbreaks when observations are typically scarce. Taking Influenza, Measles, and Dengue as examples (see dataset details in Appendix A.4), the observed performance gap in Figure 1 between a simple ARIMA model and more complex deep learning models, especially Transformer-based models, highlights a key limitation: more complex models often struggle in epidemic forecasting when prior knowledge is unavailable and observational data is scarce. Nevertheless, given the extensive and diverse outbreak data collected over decades and across various geographies, pre-training on such a broad dataset could potentially enable the development of more generalizable models with greater applicability and adaptability across different pathogens and contexts. This raises an important question: Can we leverage lessons from diverse historical disease time series to enhance epidemic forecasting accuracy for more complex transformer-based models?
To address the above question, we draw inspiration from the success of large pre-trained transformer-based models (Zhao et al., 2023) and develop a pre-trained epidemic forecasting model using extensive disease time series data to distill generalizable knowledge across pathogens and contexts. The pre-trained model can be subsequently fine-tuned for specific diseases or geographical regions. While it is possible to adapt general time series foundation models (Liang et al., 2024; Ma et al., 2024) to epidemic forecasting, their pre-trained corpus mostly consists of non-epidemic data, which may not accurately capture epidemic dynamics and infection trajectories, potentially degrading forecasting accuracy.
Furthermore, although an early effort has been made in epidemic pre-training (Kamarthi & Prakash, 2023), it overlooks critical external factors, which we refer to as environment throughout this paper. These factors such as temperature, humidity, public health policies, and interventions are well-known to influence the dynamics of disease spread in space and time (Lau et al., 2020b; Chen & Hsieh, 2012). As the environment influences both historical infection trajectories and future disease spread, it can be seen as an important confounder from the perspective of causal inference (Zhou et al., 2023; Jiao et al., 2024), which hinders the learning of true causal relations between the historical infections and future outbreaks, as shown in the structured causal model in Figure 2. Driven by the changes in the environment, a major obstacle arises as the shift in the distribution of infection trajectories between training and test datasets  (Wu et al., 2024). Insufficient consideration of such distribution shifts can obscure the relationship between historical infection data and future predictions (for a detailed discussion, see Appendix A.15), compromising a model’s ability to make accurate forecasts. As such, it is crucial to disentangle the influence of changing environments from other more intrinsic factors (e.g., a pathogen’s infection rate) affecting disease transmission dynamics. Yet, exact and explicit mechanisms by which the environment influences the disease dynamics of a particular pathogen are often not fully understood or observed, which necessitates an implicit modeling approach to identify and separate these latent environmental influences.
Figure 1: Performance gap between the ARIMA model and various non-pretrained deep learning models across selected datasets and horizons. Positive values indicate that the deep learning models underperform relative to ARIMA, highlighting the limitations of complex architectures without pretraining or prior knowledge.
Figure 2: Structural causal model for epidemic forecasting, where $Z$ refers to the environment as a confounder. $X_{sp}$ and $X_{ca}$ are the spurious and causal factors of the input.
Our Solution. We propose Covariate-Adjusted Pretraining for Epidemic forecasting (CAPE) to capture universal, environment-invariant disease patterns from historical data (Figure 3). CAPE tackles data scarcity and distribution shift by integrating pre-training with implicit environment modeling, grounded in causal analysis. Our contributions include:
  1. As one of the first epidemic pre-trained models, we propose a generalizable pre-training framework that adjusts for covariates in epidemic time series data, disentangling intrinsic disease dynamics from environmental influences and yielding environment-invariant representations.
  2. We create a comprehensive testbed for evaluating pre-trained epidemic models by assembling epidemic time series datasets across various diseases and regions. This enables thorough assessment across various scenarios, including few-shot, zero-shot, cross-location, and cross-disease settings.
  3. We demonstrate the potential of epidemic pre-training, showcasing superior performance across downstream datasets and settings. Notably, CAPE surpasses the best baseline by an average of 9.9% in the full-shot setting and 18.1% in the zero-shot setting across all tested downstream datasets.
  4. We provide an in-depth analysis of how pre-training and environment modeling affect downstream performance and mitigate the impact of distribution shifts.
2 Related Work and Problem Definition
Epidemic Forecasting Models. Traditionally, epidemic forecasting employs models like ARIMA (Sahai et al., 2020), SEIR (He et al., 2020), and VAR (Shang et al., 2021). ARIMA predicts infections by analyzing past data and errors, SEIR models population transitions using differential equations, and VAR captures linear inter-dependencies by modeling each variable based on past values. Recently, deep learning models—categorized into RNN-based, MLP-based, and transformer-based—have surpassed these methods. RNN-based models like LSTM (Wang et al., 2020) and GRU (Natarajan et al., 2023) use gating mechanisms to manage information flow. MLP-based models use linear layers (Zeng et al., 2023) or multi-layer perceptrons (Borghi et al., 2021; Madden et al., 2024) for efficient data-to-prediction mapping. Transformer-based models (Wu et al., 2021; Zhou et al., 2021, 2022) apply self-attention to encode time series and generate predictions via a decoder. However, these models are limited as they typically utilize data from only one type of disease without considering valuable insights from diverse disease datasets.
Pre-trained Time Series Models. To enable few-shot or zero-shot capabilities, transformer-based models often employ pre-training on large datasets, which typically use masked data reconstruction (Zerveas et al., 2021; Rasul et al., 2023) or promote alignment across different contexts (Fraikin et al., 2023; Zhang et al., 2022; Yue et al., 2022). For example, PatchTST (Nie et al., 2022) segments time series into patches, masks some, and reconstructs the masked segments. Larger foundational models like MOMENT (Goswami et al., 2024) aim to excel in multiple tasks (e.g., forecasting, imputation, classification) but require substantial data and computational resources. In epidemic contexts, Kamarthi et al. (Kamarthi & Prakash, 2023) pre-train a model on various diseases, improving downstream performance and highlighting pre-training’s potential in epidemic forecasting. However, these works overlook environmental influence and zero-shot ability in epidemic forecasting and lack a deep analysis of how pre-training materials impact downstream performance. In this study, we introduce environment modeling and conduct a thorough analysis of these questions (see A.5 for more discussions).
Problem Definition. Given a historical time series input: $\mathbf{x}\in\mathbb{R}^{T\times 1}$, where $T$ is the size of lookback window, the goal of epidemic forecasting is to map $\mathbf{x}$ into target trajectories (e.g. infection rates): $\mathbf{y}\in\mathbb{R}^{h}$, where $h$ denotes the size of the forecast horizon. We define $X$ and $Y$ as the random variables of input and target, respectively. During pre-training, a representation function $g_{\theta}:\mathbb{R}^{T\times 1}\to\mathbb{R}^{T\times d}$, where $d$ denotes the dimension of the latent space and $\theta$ being the parameter of the model, extracts universal properties from a large collection of epidemic time series datasets $\mathcal{D}_{\text{pre}}=\{D_{1}^{\prime},D_{2}^{\prime},\dots,D_{S}^{\prime}\}$. Then, a set of self-supervised tasks $\mathcal{T}_{\text{pre}}=\{\mathcal{T}_{i}\}_{i=1}^{R}$ is defined, and each $\mathcal{T}_{i}$ transforms a sample $\mathbf{x}\sim\mathcal{D}_{\text{pre}}$ into a new input-label pair: $(\mathbf{\tilde{x}},\mathbf{\tilde{y}})$, and optimizes a loss $\mathcal{L}_{\mathcal{T}_{i}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{pre}}}[\ell_{\mathcal{T}_{i}}(h_{\psi}(g_{\theta}(\mathbf{\tilde{x}})),\mathbf{\tilde{y}})]$, with $\ell_{\mathcal{T}_{i}}$ being the task-specific metric and $h_{\psi}$ the task-specific head.
3 Proposed Method
Pre-training models for epidemic forecasting using data from diverse contexts presents two primary challenges. First, the inherent heterogeneity stemming from varied environmental conditions (e.g., public health interventions, seasonal factors, population behaviors) necessitates a model architecture capable of accounting for these influences. Second, the pre-training objectives must be designed to foster the learning of robust representations that capture these intrinsic dynamics and generalize effectively across diverse epidemiological scenarios, including novel or data-scarce situations. We will elaborate on our solutions in the following subsections to address these challenges.
3.1 Causal Analysis for Epidemic Forecasting
In epidemic forecasting, environmental factors, such as seasonal changes and population mobility, are known to confound the relationship between historical infection patterns and future disease spread. To address this critical challenge in understanding true transmission dynamics, we draw inspiration from causal inference (Zhou et al., 2023; Jiao et al., 2024) and introduce a Structural Causal Model (Figure 2) where we treat the environment $Z$ as an unobserved confounder that influences both the independent variable (e.g., historical infections $X$) and the dependent variable (e.g., future infections $Y$). Furthermore, based on the causal decomposition assumption (Mao et al., 2022), we know $X$ is composed of (1) a spurious factor $X_{sp}$ that is environment-dependent, and (2) a causal factor $X_{ca}$ representing the underlying, more intrinsic disease transmission dynamics (e.g., pathogen infectivity) that remains environment-independent. While $X$ as a whole is observed, only $X_{ca}$ has a causal relation with target $Y$. Here, we assume that a finite set of critical factors drives the epidemic dynamics:
Assumption 3.1. The environment variable $Z$ follows a categorical distribution $p(Z)$ and takes on one of $K$ discrete environmental states, denoted as $z_{k}$. Each state $z_{k}$ is associated with a unique latent representation $\mathbf{e}_{k}\in\mathbb{R}^{h_{e}}$, capturing the unique features specific to that environment.
In constructing a predictive model for input $\mathbf{x}$, we define $\hat{Y}$ as the predicted time series $\hat{\mathbf{y}}$ and model the predictive distribution $p_{\Theta}(\hat{Y}|X)$ using $f_{\Theta}(\mathbf{x})=h_{\psi}(g_{\theta}(\mathbf{x}))$, where $\Theta=\{\theta,\psi\}$. Training typically involves maximizing the log-likelihood of $p_{\Theta}(\hat{Y}|X)$, which in practice translates to minimizing the errors over the pre-training dataset $\mathcal{D}_{\text{pre}}$: $\Theta^{*}=\arg\min_{\Theta}-\frac{1}{|\mathcal{D}_{\text{pre}}|}\sum\nolimits_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{\text{pre}}}\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}$. As the environment $Z$ impacts the distribution of the observed data through $p(X,Y|Z)=p(X|Z)p(Y|X,Z)$, we formulate the following objective:
$\begin{split}\Theta^{*}=&\arg\min\nolimits_{\Theta}-\mathbb{E}_{p(Z)}[\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p(Y,X|Z)}[\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}]].\end{split}$ (1)
The above equation suggests that the optimal $\Theta^{*}$ depends on the environment distribution $p(Z)$. If we simply maximize the likelihood $p_{\Theta}(\hat{Y}|X)$, the confounding effect of $Z$ on $X$ and $Y$ will mislead the model to capture the shortcut predictive relation between the input and the target trajectories, which necessitates implicit modeling of the environment during pre-training. Given that input infection trajectories inherently reflect the influence of the environment, it is crucial to develop mechanisms that disentangle the correlations between infection trajectories and environmental factors.
[…]
$p(Y|do(X))=\sum\nolimits_{Z}p(Y|X,Z=z)p(Z=z).$ (2)
However, obtaining detailed environmental information, or ${\bf e}_{k}$, can be challenging due to variability in data availability and quality. To address this, we resort to a data-driven approach that treats ${\bf e}_{k}$ as learnable parameters and thus allows us to dynamically infer the environmental distribution directly from the observed data. Specifically, we implement an environment estimator $q_{\phi}(Z|X)$ that infers the probability of environment states based on historical inputs together with the latent representations of each state. Then, a variational lower bound is derived (Wu et al., 2024) (see Appendix A.3):
$\log{p_{\Theta}(\hat{Y}|do(X))}\geq$ $\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]$ (3)
$-\operatorname{KL}\left(q_{\phi}(Z|X)\parallel p(Z)\right),$
where the first term maximizes the model’s predictive power and the second term regularizes the environment estimator to output a distribution close to the prior distribution $p(Z)$.
3.2 Model Instantiation
Figure 3: (a) CAPE encoder and environment estimator with latent representations; (b) Hierarchical environment contrasting for temporal and environment representations; (c) Random masking and reconstruction with environment estimation to capture universal patterns; (d) EM algorithm to iteratively optimize model parameters and environment representations.
To instantiate and train a model that performs the covariate adjustment, we need to model the environment estimator $q_{\phi}(Z|X)$ and the predictor $p_{\Theta}(\hat{Y}|X,Z)$.
Latent Environment Estimator $q_{\phi}(Z|X)$. We approximate the distribution of environmental contexts $p(Z)$ using a latent environment estimator $q_{\phi}(Z|X)$. Epidemic infection trajectories are profoundly shaped by dynamic environmental influences—such as the implementation and lifting of non-pharmaceutical interventions (NPIs), seasonal drivers affecting pathogen viability (e.g., influenza seasonality), or shifts in population behavior—which can vary significantly over different periods of an outbreak. Modeling the impact of these potentially complex and time-varying influences across an entire long infection trajectory can be inefficient and may obscure shorter-term but critical environmental shifts. Therefore, we apply patching (Nie et al., 2022) to manage the granularity in our environment estimation. This approach allows the model to identify distinct environmental ”regimes” or periods within an epidemic wave that might correspond to different dominant external factors (e.g., a period of strict lockdown versus a period of reopening). Such granularity helps prevent overly coarse or overly fine-grained estimations that could obscure epidemiologically relevant shifts in transmission dynamics, such as changes in the effective reproduction number or the shape of the epidemic curve. We divide the historical infection input $\mathbf{x}$ (e.g., weekly case counts) into $C$ non-overlapping temporal patches, $\mathbf{x}=[\mathbf{x}_{1},\dots,\mathbf{x}_{C}]$, where $\mathbf{x}_{c}\in\mathbb{R}^{T/C}$. A self-attention layer $f_{\text{enc}}$ then captures temporal dependencies between these patches, producing contextualized representations of the local infection trend $\mathbf{h}^{(l)}_{c}=f_{\text{enc}}(\mathbf{x}^{(l)}_{c})$ for each patch at layer $l$. Subsequently, the environment state and the patch representation are projected to the same dimension and we calculate the cross-attention score using a softmax function as the environment probability:
$\pi_{k,c}^{(l)}=\operatorname{Softmax}\left((\mathbf{W}_{k}^{(l)}\mathbf{e}_{k})^{\top}\cdot(\mathbf{W}_{s}^{(l)}\mathbf{h}_{c}^{(l)})\right),$ (4)
where $\pi_{k,c}^{(l)}=q_{\phi}(Z=z_{k}|X=x_{c})$ is the output probability of the environment $z_{k}$ for the $c$-th patch, and $\mathbf{W}_{k}^{(l)}$ and $\mathbf{W}_{s}^{(l)}$ are learnable linear layers. This mechanism allows the model to not only leverage the observed infection pattern within the patch $\mathbf{h}_{c}^{(l)}$ but also to consider a set of learned environmental influences $\textbf{e}_{k}$. The $\textbf{e}_{k}$ are optimized to capture distinct types of environmental conditions, expressed as a vector of channel-wise weights, relevant to disease spread across many historical epidemics encountered during pre-training. This design is crucial as it enables the model to potentially infer the influence of combinations or intensities of environmental factors not explicitly seen together in the training data for a specific downstream task, thereby enhancing its ability to generalize to novel epidemic scenarios or evolving outbreaks.
Epidemic Predictor $p_{\Theta}(\hat{Y}|X,Z)$. Unlike previous studies, which do not explicitly model environment states, we incorporate these states directly into the patched infection trend using their latent representations $\mathbf{e}_{k}$. Specifically, we model the predictor $p_{\Theta}(\hat{Y}|X,Z)$ in Eq (2) by employing a weighted sum over the combined representations of each environment and the infection trend using Hadamard product, i.e., $f_{\text{enc}}(\mathbf{x}_{c}^{(l)})\odot\mathbf{e}_{k}$. Finally, we apply a feed-forward layer is applied. This step gives us the adjusted environment-invariant representation for $X_{ca}$ in Figure 2, and we conduct such adjustment in a layer-wise manner, which can be expressed as:
$\small\mathbf{x}^{(l+1)}_{c}=\sigma\left(\mathbf{W}_{f}^{(l)}\sum\nolimits_{k=1}^{K}\pi_{k,c}^{(l)}\left[f_{enc}({\mathbf{x}_{c}^{(l)}})\odot\mathbf{e}_{k}\right]\right),$ (5)
where $\sigma$ represents the activation function and $\mathbf{W}_{f}^{(l)}$ denotes the learnable parameters of the feedforward layer. Assuming $L$ layers are stacked, we acquire the final representation $\mathbf{x}^{(L)}=[\mathbf{x}^{(L)}_{1},\mathbf{x}^{(L)}_{2},\ldots\mathbf{x}^{(L)}_{C}]=g_{\theta}({\bf x})\in\mathbb{R}^{C\cdot d}$ and apply a task-specific head to predict the target variable $\mathbf{y}=h_{\psi}(\mathbf{x}^{(L)})$, where $h_{\psi}$ is a linear transformation.
3.3 Pre-training Objectives for Epidemic Forecasting
CAPE captures diverse epidemic time series dynamics through self-supervised learning tasks that identify universal patterns in the pre-training dataset. While previous studies neglected the confounding effects of environmental factors on input-label pairs in $\mathcal{T}_{\text{pre}}$, CAPE seamlessly integrates environment estimation into the self-supervised framework.
Random Masking with Environment Estimation. To learn from large unlabeled epidemic time series, we adopt a masked modeling task (Kamarthi & Prakash, 2023; Goswami et al., 2024), masking 30% of input patches (Figure 3(c)). As shown in Figure 2, input $X$ is influenced by the environment $Z$, so reconstructing patches requires modeling both temporal and environmental dependencies. Unlike previous methods that ignore this, we use backdoor adjustment via an environment estimator $q_{\phi}(Z|X)$ to obtain environment-invariant representations. During pre-training, the input $\mathbf{x}$ is masked to produce pairs $(\tilde{\mathbf{x}},\mathbf{x})$, with the reconstruction $\hat{\mathbf{x}}=h_{\psi}(g_{\theta}(\tilde{\mathbf{x}}))$ trained using Mean Squared Error: $\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})=\text{MSE}(\hat{\mathbf{x}},\mathbf{x})$.
Hierarchical Environment Contrasting for Robust Epidemiological Contextualization. During optimization, inconsistent environment representations can occur if the model, when processing two infection trajectories (${\bf x}$ and ${\bf x}^{\prime}$) that share an identical overlapping time period, estimates different environmental contexts for that common period based on the distinct temporal context. To address the challenge and improve consistency in representing environments for adjacent epidemic time series segments, we propose Hierarchical Environment Contrasting. Specifically, We define an aggregated environment representation as the confluence of environmental factors, $\hat{\mathbf{e}}^{(l)}_{c}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k,c}^{(l)}$, for each patch $c$, allowing it to reflect a nuanced, weighted combination of $K$ learned latent environments ($\mathbf{e}_{k}$) rather than a single discrete state. Our contrastive feature, $\hat{\mathbf{E}}_{j,c}^{(l)}=\sigma(\mathbf{W}_{f}^{(l)}(\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}))$, is designed to capture this essential interaction. It explicitly combines the aggregated environment representation ($\hat{\mathbf{e}}^{(l)}_{c}$) with the local infection trend representation ($\mathbf{h}^{(l)}_{c}$) via a Hadamard product ($\odot$), enabling the model to learn how specific environmental profiles alter observed infection trends (e.g., by changing growth rates). The overarching goal, realized through the contrastive loss (Equation 3.3), is to ensure these learned features $\hat{\mathbf{E}}_{j,c}^{(l)}$ are consistent across genuinely comparable epidemiological contexts and distinct for dissimilar underlying situations, even if superficial case counts appear similar, fostering robust, discriminative representations of environmental influence. The patch-wise contrastive loss is:
$\mathcal{L}_{\text{CL}}(j,c)=-\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,c)}$
$\hskip 28.45274pt+\log\left(\sum\nolimits_{b\in B}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(b,c)}\right)+\mathbb{I}_{j\neq b}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(b,c)}\right)\right)$
$\hskip 28.45274pt+\log\left(\sum\nolimits_{t\in\Omega}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,t)}\right)+\mathbb{I}_{c\neq t}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(j,t)}\right)\right).$
where $B$ is the batch, $\Omega$ is the set of overlapping patches, and $\mathbb{I}$ is the indicator function. For more details, please refer to Appendix A.4.
Given a batch $\mathbf{X}\in\mathbb{R}^{B\times T}$, our pre-training loss is: $\mathcal{L}_{\text{final}}=\sum\nolimits_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}}^{\prime(L)})$, where $L$ is the number of layers and $\alpha$ balances reconstruction and contrastive losses.
3.4 Optimization of the CAPE Framework
To maximize the variational lower bound in Eq. (3), we adopt the Expectation-Maximization (EM) algorithm to iteratively update latent environments and the epidemic predictor (see Appendix A.5 for the pseudo algorithm).
E-Step: Estimating Latent Environments. We estimate the environment states $Z$ and their distribution $p(Z)$ by maximizing the expected likelihood of $p(Y|Z)$. Fixing the epidemic predictor $p_{\Theta}(\hat{Y}|X,Z)$ and environment estimator $q_{\phi}(Z|X)$, we treat them as oracles such that $p_{\Theta}(\hat{Y}|X,Z)=p(Y|X,Z)$ and $q_{\phi}(Z|X)=p(Z)$. The environment representations $\mathbf{E}$ are updated by minimizing the expected reconstruction loss: $\mathbf{E}^{t+1}\leftarrow\arg\min\nolimits_{\mathbf{E}}\left[\mathbb{E}_{\mathbf{x}\sim p(X)}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})\right]$. This yields the updated distribution $p^{t+1}(Z)=q^{t+1}_{\phi_{t}}(Z)$ and representations $\mathbf{E}^{t+1}$.
M-Step: Optimizing Epidemic Predictor. With $\mathbf{E}^{t+1}$ fixed, we provide the following theorem:
Theorem 3.1.
Assuming $q^{t+1}_{\phi_{t}}(Z)=p^{t+1}(Z)$ and an L2 norm is applied on $\phi$, the variational lower bound in Eq. (3) can be approximated as follows: $\mathbb{E}_{p(X)}\left[\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\left[\log p^{t+1}_{\Theta_{t+1}}(\hat{Y}|X,Z)\right]\right]-C$, which is equivalent to minimizing the expected reconstruction loss $\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{recon}(\mathbf{x},\hat{\mathbf{x}})]$
Further theoretical details are provided in Appendix A.3. Therefore, we optimize model parameters by minimizing the expected reconstruction loss and incorporating a contrastive loss to regularize the environment estimator: $\Theta_{t+1}\leftarrow\arg\min\nolimits_{\Theta}\left[\mathcal{L}_{\text{final}}(\mathbf{X},\hat{\mathbf{X}},\mathbf{E}^{t+1})\right]$.
Table 1: Univariate forecasting results with horizons ranging from 1 to 16 future steps. The lookback window length is set to 36 and all models are evaluated using MSE. Note that performance rankings are distinguished by color coding: Best, Second Best, Third Best. $\Delta(\%)$ stands for the relative improvement of CAPE over the baselines in terms of average MSE over all horizons.
Dataset Horizon Statistical Model RNN-Based MLP-Based Transformer-Based CAPE
ARIMA LSTM GRU Dlinear Non-Pre-trained Pre-trained
Informer Autoformer Fedformer PEM MOMENT PatchTST
ILI USA 1 0.138 0.338 0.259 0.220 0.175 0.457 0.368 0.179 0.269 0.195 0.155
2 0.203 0.377 0.301 0.247 0.370 0.710 0.380 0.226 0.321 0.264 0.200
4 0.354 0.458 0.386 0.376 0.517 0.670 0.433 0.304 0.397 0.385 0.270
8 0.701 0.579 0.529 0.506 0.597 0.842 0.570 0.538 0.510 0.535 0.404
16 1.121 0.691 0.626 0.617 0.812 0.835 0.701 0.570 0.610 0.485 0.516
Avg 0.503 0.489 0.420 0.393 0.494 0.703 0.490 0.363 0.421 0.373 0.309
$\Delta$ (%) 38.57% 36.81% 26.43% 21.37% 37.45% 56.05% 36.94% 14.88% 26.60% 17.16% -
ILI Japan 1 0.358 1.426 1.213 1.016 0.405 0.515 0.525 0.470 0.325 0.413 0.290
2 0.772 1.635 1.458 1.294 0.666 0.855 1.151 0.755 0.586 0.698 0.535
4 1.720 1.975 1.870 1.758 1.234 1.150 1.455 1.207 1.082 1.147 0.944
8 2.981 2.373 2.365 2.285 1.688 1.866 2.012 1.810 1.706 1.708 1.650
16 2.572 2.023 2.010 2.007 1.551 2.654 4.027 1.766 2.054 1.688 1.911
Avg 1.680 1.886 1.783 1.672 1.109 1.408 1.834 1.202 1.151 1.131 1.066
$\Delta$ (%) 36.55% 43.48% 40.21% 36.24% 3.88% 24.29% 41.88% 11.31% 7.38% 5.74% -
Measles 1 0.071 0.182 0.143 0.133 0.066 0.203 0.321 0.085 0.113 0.094 0.083
2 0.120 0.223 0.176 0.184 0.153 0.257 0.817 0.128 0.138 0.127 0.112
4 0.225 0.310 0.258 0.296 0.288 0.331 0.226 0.213 0.186 0.205 0.161
8 0.483 0.567 0.471 0.512 0.501 0.671 0.403 0.417 0.351 0.377 0.310
16 1.052 1.110 1.013 1.088 0.904 1.115 0.754 0.806 0.818 0.722 0.752
Avg 0.390 0.478 0.412 0.443 0.382 0.515 0.504 0.330 0.321 0.305 0.269
$\Delta$ (%) 31.03% 43.72% 34.71% 39.28% 29.58% 47.77% 46.63% 18.49% 16.20% 11.80% -
Dengue 1 0.244 0.250 0.261 0.224 0.255 0.525 0.521 0.225 0.420 0.240 0.223
2 0.373 0.343 0.343 0.316 0.450 0.807 0.670 0.314 0.579 0.334 0.302
4 0.696 0.564 0.579 0.560 0.798 0.957 0.766 0.571 0.661 0.586 0.561
8 1.732 1.168 1.183 1.256 1.239 1.684 1.539 1.223 1.308 1.292 1.046
16 4.082 3.876 3.315 3.109 2.659 3.364 2.934 3.376 2.532 2.537 2.509
Avg 1.426 1.240 1.136 1.093 1.080 1.467 1.286 1.142 1.100 1.000 0.892
$\Delta$ (%) 37.45% 28.06% 21.48% 18.39% 17.41% 39.20% 30.64% 21.89% 18.91% 10.80% -
Covid 1 33.780 22.592 22.009 23.811 34.161 42.049 28.130 25.088 32.376 23.645 21.548
2 33.193 23.460 22.542 24.809 24.883 30.631 28.059 23.123 35.418 25.047 22.224
4 32.482 24.729 24.816 26.345 31.328 41.029 29.432 23.889 36.251 24.224 22.476
8 36.573 31.019 33.934 33.081 35.964 55.812 41.791 31.217 40.429 31.548 28.403
16 42.910 43.820 41.432 47.561 50.244 47.993 69.976 51.265 52.590 43.309 40.555
Avg 35.787 29.124 28.947 31.121 35.316 43.503 39.478 30.917 39.413 29.555 26.559
$\Delta$ (%) 25.79% 8.81% 8.25% 14.66% 24.80% 38.95% 32.72% 14.10% 32.61% 10.14% -
Table 2: Few-shot learning results with horizons ranging from 1 to 16 future steps. The length of the lookback window is set to 36. Each model is evaluated after being trained on 20%, 40%, 60%, and 80% of the full training data. $\Delta(\%)$ stands for the relative improvement after training with 20% more data in terms of average MSE over all horizons. The full result is shown in Appendix A.9.
Dataset CAPE PatchTST Dlinear MOMENT PEM
20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100%
ILI USA 2.121 1.400 0.760 0.369 0.309 2.114 1.219 0.677 0.401 0.373 2.822 1.594 0.816 0.412 0.346 3.990 1.847 0.913 0.459 0.381 2.143 1.261 0.681 0.419 0.353
$\Delta$(%) - 33.99% 45.71% 51.45% 16.26% - 42.34% 44.45% 40.77% 6.98% - 43.53% 48.78% 49.51% 16.02% - 53.69% 50.58% 49.72% 17.00% - 41.13% 46.00% 38.33% 15.76%
Dengue 13.335 6.386 2.356 1.511 0.892 13.712 7.304 2.771 1.678 0.984 15.828 8.420 2.850 1.748 1.080 15.697 7.536 2.816 1.733 1.358 12.90 7.055 2.745 1.707 0.964
$\Delta$(%) - 52.07% 63.12% 35.87% 40.95% - 46.72% 62.06% 39.43% 41.39% - 46.81% 66.15% 38.64% 38.19% - 52.00% 62.63% 38.45% 21.65% - 45.32% 61.09% 37.79% 43.51%
Measles 0.483 0.600 0.381 0.285 0.269 0.863 0.834 0.448 0.359 0.306 1.194 1.130 0.602 0.478 0.394 1.661 0.915 0.425 0.471 0.500 0.670 0.896 0.430 0.364 0.306
$\Delta$(%) - -24.22% 36.50% 25.20% 5.61% - 3.36% 46.25% 19.91% 14.81% - 5.36% 46.64% 20.63% 17.58% - 44.91% 53.55% -10.59% -6.16% - -33.87% 51.91% 15.35% 15.93%
4 Experiment
4.1 Experiment Setup
Datasets and Baselines. For pre-training CAPE, PatchTST, and PEM, we manually collected 17 distinct weekly-sampled diseases from Project Tycho (van Panhuis et al., 2018). For evaluation, we utilize five downstream datasets covering various diseases and locations: ILI USA (Centers for Disease Control and Prevention, 2023a), ILI Japan (National Institute of Infectious Diseases, 2023), COVID-19 USA (Dong et al., 2020), Measles England (Lau et al., 2020a), and Dengue across countries (OpenDengue, 2023). Additionally, RSV (Centers for Disease Control and Prevention, 2023c) and MPox (Centers for Disease Control and Prevention, 2023b) infections in the US are used to test zero-shot performance. More details can be found in Appendix A.4. We adopt baseline models from the comprehensive EpiLearn toolkit (Liu et al., 2024a), comparing CAPE against two categories: non-pretrained and pre-trained models. Non-pretrained baselines include classical statistical models such as ARIMA (Panagopoulos et al., 2021), configured in an online forecasting setting (Pham et al., 2022; Abdulmajeed et al., 2020) where parameters are updated with each new sample. We also consider deep learning models, including RNN-based approaches like LSTM and GRU (Wang et al., 2020; Natarajan et al., 2023), the linear model DLinear (Zeng et al., 2023), and transformer-based architectures (Wu et al., 2021; Zhou et al., 2021, 2022). Pre-trained baselines include state-of-the-art models such as PatchTST (Nie et al., 2022), PEM (Kamarthi & Prakash, 2023), and the time series foundation model MOMENT (Goswami et al., 2024). Further experimental details are provided in Appendix A.5.
4.2 Fine-Tuning (Full-Shot Setting)
For non-pre-trained models, we train the entire model on the training split, while for pre-trained models, we fine-tune on downstream datasets by transferring the task-specific head $h_{\psi}$ from pre-training to the forecasting task. We evaluate short-term and long-term performance by reporting MSE across horizons from 1 to 16. From Table 1, we observe: (a) CAPE achieves the best average MSE across all downstream datasets. It outperforms the best baseline by 9.91% on average and up to 14.85%. On the COVID dataset, CAPE performs best across all horizons, showing effectiveness on novel diseases. (b) Models like PEM, PatchTST, and MOMENT consistently rank second or third on 4 out of 5 downstream datasets. The best pre-trained model (excluding CAPE) outperforms the best non-pre-trained model by 6.223% on average. Among them, PatchTST has the highest average performance, surpassing PEM by 5.51% and MOMENT by 10.45%. Additionally, PEM outperforms MOMENT by 4.86%, indicating the importance of epidemic-specific pre-training. (c) Informer consistently outperforms Autoformer and Fedformer by 24.40% and 17.90% respectively, due to its sparse attention mechanism that reduces overfitting. Informer also surpasses Dlinear by 1.90%, suggesting that careful selection of model size and parameters is crucial for optimal performance. (d) Furthermore, environment modeling proves valuable, as CAPE consistently outperforms PatchTST, which shares a similar design. While both models are pre-trained on the epidemic-specific datasets, CAPE surpasses PatchTST by 11.13%.
4.3 Few-Shot Performance, Zero-Shot Performance, and Ablation Study
Few-Shot Forecasting. In real-world scenarios, predicting outbreaks of diseases unknown or in new locations is challenging for purely data-driven models due to limited initial data. Thus, few-shot or zero-shot forecasting capabilities are essential for epidemic models. To simulate a few-shot scenario, we reduce the original training data from 100% to [20%, 40%, 60%, 80%]. We report the average MSE across 1 to 16 time steps. From Table 11, we make the following observations: (a) With an increasing volume of training materials, the performance of all models consistently improves. (b) CAPE achieves the best performance in most scenarios, demonstrating the superior few-shot ability. (c) Compared with models pre-trained on epidemic-specific datasets, Dlinear failed to achieve better performance when only 20% of training data is available. However, Dlinear is able to outperform MOMENT on ILI USA and Measles datasets when both models are trained or fine-tuned using 20% training data, which indicates the importance of pre-training. (d) Though CAPE achieves the best average performance on the ILI USA dataset when the training material is reduced, it achieves a good performance in short-term forecasting from 1 to 4 weeks (see Appendix A.9).
Table 3: Zero-shot performance with a lookback window length of 12. All results are averaged over 4 weeks or days in the future. $\Delta(\%)$ stands for the relative improvement over the baselines.
Dataset $\Delta$ (%) CAPE PatchTST PEM MOMENT
ILI USA 9.26% 0.147 0.164 0.162 0.549
ILI Japan 17.06% 0.705 0.907 0.850 2.062
Measles 3.97% 0.145 0.167 0.159 0.533
MPox 20.00% 0.0004 0.0005 0.0005 0.0013
Dengue (mixed) 10.17% 0.371 0.427 0.413 1.624
RSV 26.06% 0.834 1.128 1.260 1.849
Covid (daily interval) 13.80% 5.173 6.001 6.320 18.881
Table 4: Ablation study.
Dataset Model H=1 H=2 H=4 H=8 H=16 Avg
ILI USA CAPE 0.155 0.200 0.270 0.404 0.516 0.309
w/o Env 0.326 0.448 0.508 0.642 0.735 0.532
w/o Contrast 0.174 0.241 0.335 0.492 0.570 0.363
w/o Pretrain 0.158 0.202 0.283 0.408 0.545 0.319
Measles CAPE 0.069 0.096 0.155 0.280 0.743 0.269
w/o Env 0.083 0.111 0.168 0.407 0.755 0.304
w/o Contrast 0.090 0.124 0.276 0.431 0.801 0.344
w/o Pretrain 0.074 0.113 0.223 0.402 0.816 0.326
Dengue CAPE 0.218 0.301 0.540 1.193 2.210 0.892
w/o Env 0.232 0.316 0.484 1.089 3.622 1.149
w/o Contrast 0.198 0.273 0.460 1.128 3.329 1.078
w/o Pretrain 0.210 0.276 0.449 1.115 3.759 1.162
Zero-Shot Forecasting. To further demonstrate the potential of our model, we evaluate CAPE in a zero-shot setting. Specifically, for transformer-based models, we retain the pre-training head and freeze all parameters during testing. All models are provided with a short input sequence of 12 time steps and tasked with predicting infections for the next 4 time steps. From Table 4, we make the following observations: (a) CAPE outperforms baselines across all downstream datasets, showing superior zero-shot forecasting ability. (b) Models pre-trained on epidemic-specific datasets achieve better performance compared to those pre-trained without epidemic-specific data (MOMENT). This indicates the necessity of choosing domain-specific materials for pre-training.
Ablation Study. We conducted an ablation study to assess CAPE’s components (Table 4). Replacing environment estimators with non-disentangling self-attention layers consistently worsened performance across all datasets, notably increasing ILI USA’s MSE from 0.309 to 0.532. Similarly, removing contrastive loss while retaining environment estimators raised Measles’ MSE from 0.269 to 0.344, with smaller increases for ILI USA and Dengue. Training CAPE directly on downstream datasets without pre-training also decreased performance, with MSE rising to 0.319 (ILI USA), 0.326 (Measles), and 1.162 (Dengue), though less than removing environment estimation. These results indicate that all CAPE components are essential for optimal forecasting.
4.4 Transferability
Cross-Location: We include measles data from the USA in the pre-training dataset. To evaluate our model’s ability to adapt to cross-region data, we incorporate measles outbreak data from the UK into the downstream datasets. As shown in Table 4, the pre-trained CAPE outperforms the non-pre-trained version by 17.48%. While we pre-train our model with influenza data from the USA, the zero-shot evaluation on the influenza outbreak in Japan also shows superior performance, underscoring the crucial role of pre-training in enabling generalization to novel regions. Cross-Disease: While we include various types of diseases in our pre-training dataset, novel diseases including Dengue (non-respiratory) and COVID-19 that are unseen in the pre-training stage are incorporated during the downstream evaluation. The ability of our model to adapt to novel diseases is proven compared to the version not pre-trained on the Dengue dataset (Table 4), improving which by 23.24%, as well as the superior zero-shot performance on the COVID dataset (Table 4), which surpasses the MOMENT that is not pre-trained on other diseases by 72.60%. Cross-Interval: While we only pre-train using weekly-sampled data, our model outperformed the non-pre-trained version on the irregularly sampled Dengue dataset and on the daily-sampled COVID-19 dataset, demonstrating robustness to different time intervals.
4.5 Deeper Analysis
Impact of Pre-Training Epochs: Evaluating four downstream datasets (Figure 5), we find that increasing pre-training epochs consistently improves performance on Measles and COVID datasets but degrades it for ILI USA. Additionally, models with more environment states $K$ perform better as pre-training epochs increase. Impact of Pre-Training Materials: We examine potential biases in our pre-training dataset by splitting it into respiratory and non-respiratory diseases. As shown in Figure 4, with similar volumes of pre-train data, the model performs better when the tested disease types align with the pre-train data. However, the size of the pre-train material has a stronger impact. Impact of Pre-Training Material Scale: To explore how the pre-training material scale affects downstream performance, we scaled the original pre-training dataset and test on downstream datasets. As shown in Figure 6, a sudden performance boost is observed at around a 60% reduction for both Measles and Dengue datasets.
Figure 4: Downstream performance when the model is pre-trained with either respiratory or non-respiratory data.
Tackling Distribution Shift: We define distribution shifts as changes in infection patterns from training to test data. We compute the Central Moment Discrepancy (CMD)(Zellinger et al., 2017) between training and test distributions for each disease. As shown in Figure 7, CAPE with environment estimation achieves the lowest CMD scores, highlighting its effectiveness in mitigating distribution shifts. Disentangling Disease Dynamics: We assess CAPE’s ability to capture intrinsic disease dynamics by extracting latent embeddings and computing the Davies-Bouldin Index (DBI) across different dataset pairs. As shown in A.13, CAPE consistently yields lower DBI scores than PatchTST, indicating better separation of disease-specific patterns from environmental influences.
Figure 5: Downstream performance vs. environments distribution. More results in A.10.
Figure 6: Downstream performance across pre-training ratios. More datasets are in  A.11.
Figure 7: CMD scores w/ and w/o environment estimate between train/test sets. (A.12)
5 Conclusion
We present CAPE, a covariate-adjusted pre-training framework for epidemic forecasting that captures environment-invariant disease dynamics by combining pre-training with implicit environment modeling grounded in causal analysis. By learning latent environments and applying backdoor adjustment, CAPE reduces distribution shifts. Extensive experiments across diseases, regions, and scenarios demonstrate strong generalization, with consistent gains in few-shot and zero-shot settings. Future directions can include incorporating uncertainty measurement, scaling to larger datasets, and providing interpretable environment representations.
Appendix A Appendix
A.1 Potential Impact
This paper advances the interdisciplinary fields of machine learning and epidemiology by enhancing the accuracy of epidemic forecasting in data-limited settings. We outline the potential impacts as follows:
Early Insights. We provide novel insights into how pre-training and environment modeling improve epidemic forecasting. Our results demonstrate that pre-training significantly enhances model accuracy, with gains increasing as more pre-training data is incorporated. This finding paves the way for future research to develop foundational models in epidemic analysis using larger datasets. Additionally, we confirm the importance of accounting for both inherent disease dynamics and environmental factors to achieve robust forecasting performance.
Social Impact. Epidemic time series data are often sparse due to limited sampling rates, hindering public health organizations’ ability to accurately predict infections during novel disease outbreaks. This paper addresses this challenge by showcasing the few-shot and zero-shot forecasting capabilities of pre-trained models. These capabilities can provide powerful tools for early warning and timely intervention, ultimately supporting more effective public health responses and safeguarding communities against emerging infectious diseases.
A.2 Limitations
While CAPE shows strong performance, it has several limitations. First, it does not currently incorporate uncertainty estimation, which is important for risk-aware decision-making. Second, its effectiveness may be constrained by the scale and diversity of available data. Lastly, the learned environment representations lack interpretability, limiting their transparency and potential for insight in public health contexts.
Though not explicitly designed for uncertainty, CAPE supports Dropout Injection (Ledda et al., 2023) at inference. We use 25 stochastic forward passes on the ILI USA dataset to compute 95% confidence intervals:
Table 5: CAPE Forecasting Results with 95% Confidence Intervals
Horizon CAPE 95% CI
1 $0.174\pm 0.0036$
2 $0.235\pm 0.0016$
4 $0.293\pm 0.0012$
8 $0.433\pm 0.0010$
A.3 Theoretical Analysis
A.3.1 Derivation for do-operation
We derive a tractable form for $\log p_{\Theta}(\hat{Y}|do(X))$, leveraging two rules of do-calculus.
Do-Calculus Rules.
Consider a causal DAG $\mathcal{A}$ with nodes $B$, $D$, and $Z$. Let $\mathcal{A}_{B}$ denote the intervened graph by removing all arrows entering $B$, and $\mathcal{A}^{B}$ the graph by removing all arrows leaving $B$. The rules are:
1. Action/Observation Exchange:
$P(D|do(B),do(Z))=P(D|do(B),Z),$
if $(D\perp\!\!\!\perp Z|B)$ in $\mathcal{A}^{Z}_{B}$.
2. Insertion/Deletion of Actions:
$P(D|do(B),do(Z))=P(D|do(B)),$
if $(D\perp\!\!\!\perp Z|B)$ in $\mathcal{A}^{BZ}$.
We consider a causal graph with variables $Z$, $X$, and $\hat{Y}$, as shown in Figure 2. Starting with the law of total probability:
[…]
Step 1: Action/Observation Exchange. Using $(\hat{Y}\perp\!\!\!\perp X|Z)$ in $\mathcal{A}_{X}$, we apply the exchange rule:
$P(\hat{Y}|do(X),Z=z)=P(\hat{Y}|X,Z=z).$ (7)
Step 2: Insertion/Deletion of Actions. Using $(Z\perp\!\!\!\perp X)$ in $\mathcal{A}_{X}$, we simplify:
$P(Z=z|do(X))=P(Z=z).$ (8)
Substituting these into (6), we obtain:
[…]
The result can be compactly written as:
$p_{\Theta}(\hat{Y}|do(X))=\mathbb{E}_{p_{0}(Z)}\left[p_{\theta}(\hat{Y}|X,Z)\right],$ (10)
where $p_{0}(Z)$ denotes the prior distribution of environments. Do-calculus rules simplify interventional distributions by leveraging independence properties, enabling tractable objectives for causal inference.
A.3.2 Derivation for Variational Lower Bound
Below we show the derivation for the variational lower bound in Eq. (3).
$\quad\log\sum_{z}p_{\Theta}(\hat{Y}|X,Z=z)P(Z=z)$ (11)
$=\log\sum_{z}p_{\Theta}(\hat{Y}|X,Z=z)p(Z=z)\frac{q_{\phi}(Z=z|X)}{q_{\phi}(Z=z|X)}$
$\geq\sum_{z}q^{t+1}_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)\frac{p(Z=z)}{q_{\phi}(Z=z|X)}\quad$
(Jensen’s Inequality)
$=\sum_{z}q_{\phi}(Z=z|X)\log p_{\Theta}(\hat{Y}|X,Z=z)-\sum_{z}q_{\phi}(Z=z|X)\log\frac{q_{\phi}(Z=z|X)}{p(Z=z)}$
$=\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]-KL\left(q_{\phi}(Z|X)\parallel p(Z)\right)$
(a) Previous Method
(b) Ours
Figure 8: Comparison between previous approach and our implementation.
A.3.3 Proof of Theorem 3.1
In this section, we provide the theoretical analysis of our optimization method, and an over is shown in Figure 8. Before proving Theorem 3.1, we prove the following theorem:
Theorem A.1.
Given the environment estimator before update $q_{\phi_{t}}$ and after update $q_{\phi_{t+1}}$, assuming the training process converges, then minimizing the KL divergence from $q_{\phi_{t+1}}$ to $q_{\phi_{t}}$, i.e., $\text{KL}(q_{\phi_{t+1}}||q_{\phi_{t}})$, is equivalent to applying an L2 norm on the parameters of the environment estimator $q_{\phi}$.
Proof.
Since $q_{\phi}$ only involves linear transformations and a softmax function, we argue that minimizing the KL loss between $q_{\phi}^{t}$ and $q_{\phi}^{t+1}$ is equivalent to minimizing the difference between $\phi^{t}$ and $\phi^{t+1}$.
Firstly, the logits $s_{i}$ are computed as: $s_{i}=\mathbf{W}_{k}^{(l)}\mathbf{e}^{k}$. A change in $\mathbf{W}_{k}^{(l)}$ , denoted as $\Delta\mathbf{W}_{k}^{(l)}$ , modifies $s_{i}$ as $\Delta s_{i}=(\Delta\mathbf{W}_{k}^{(l)})\mathbf{e}^{k}$. The change in $s_{i}$ is linear with respect to $\Delta\mathbf{W}_{k}^{(l)}$ and $\mathbf{e}^{k}$ Therefore, a small change in $\mathbf{W}_{k}^{(l)}$ leads to proportionally small changes in $s_{i}$.
Secondly, the Softmax function introduces nonlinear coupling between logits $s_{i}$ , as the output probabilities $\pi_{i}$ depend not only on $s_{i}$ but also on all other logits $s_{j}$ For a small change in $s_{i}$ , the change in $\pi_{i}$ can be approximated using the gradient of the Softmax $\frac{\partial\pi_{i}}{\partial s_{i}}=\pi_{i}(1-\pi_{i})$, $\frac{\partial\pi_{i}}{\partial s_{j}}=-\pi_{i}\pi_{j}\quad\text{for }i\neq j$. Thus, a change in $\mathbf{W}k^{(l)}$ affects $\pi_{i}$ as $\Delta\pi_{i}\approx\pi_{i}(1-\pi_{i})\Delta s_{i}-\sum_{j\neq i}\pi_{i}\pi_{j}\Delta s_{j}$. This implies that the change of the output $\Delta\pi_{i}$ shrinks proportionally with the change of parameters $\Delta\mathbf{W}_{k}$.
Lastly, for small parameter changes, we can approximate the KL divergence between the two categorical distributions using a second-order Taylor expansion around $q_{\phi}^{t}$. The KL divergence is defined as:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)=\sum_{i}\pi_{i}^{t}\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right).$ (12)
Expanding $\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ around $\pi_{i}^{t}$ using the Taylor series for $\log(1+x)$ where $x=-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}$ and keeping terms up to second order, we obtain:
$\log\left(\frac{\pi_{i}^{t}}{\pi_{i}^{t+1}}\right)$ $=\log\left(1+\frac{\pi_{i}^{t}-\pi_{i}^{t+1}}{\pi_{i}^{t}}\right)=\log\left(1-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)\approx-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}.$ (13)
Substituting this into the KL divergence expression and ignoring higher-order terms (since $\Delta\pi_{i}$ is small), we get:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)$ $\approx\sum_{i}\pi_{i}^{t}\left(-\frac{\Delta\pi_{i}}{\pi_{i}^{t}}-\frac{1}{2}\left(\frac{\Delta\pi_{i}}{\pi_{i}^{t}}\right)^{2}\right)=-\sum_{i}\Delta\pi_{i}-\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$ (14)
Since $\sum_{i}\Delta\pi_{i}=0$ (as probabilities sum to one), the first term vanishes, leaving:
$\text{KL}\left(q_{\phi}^{t}\parallel q_{\phi}^{t+1}\right)\approx\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}.$ (15)
Substituting the expression for $\Delta\pi_{i}$ from above and noting that $\Delta s_{i}$ is linear in $\Delta\phi$, we observe that the KL divergence is a quadratic function of $\Delta\phi$. Therefore, for small $\Delta\phi$, the KL divergence can be approximated as:
$\mathcal{L}_{\text{KL}}\approx\frac{1}{2}\sum_{i}\frac{(\Delta\pi_{i})^{2}}{\pi_{i}^{t}}\propto\|\Delta\phi\|^{2}=\|\phi^{t}-\phi^{t+1}\|^{2}.$ (16)
Thus, minimizing the KL divergence is approximately equivalent to minimizing $\|\phi^{t}-\phi^{t+1}\|^{2}$. In our setting, we use weight decay to regularize the model, which indirectly helps control the KL loss. Weight decay adds an L2 penalty to the loss function that encourages smaller parameter values, effectively shrinking the magnitude of the weights during training. The update rule for parameters with weight decay is given by: $\phi_{t+1}=\phi_{t}-\alpha\cdot\nabla_{\phi}\mathcal{L}_{\text{original}}-\alpha\cdot\lambda\phi_{t}$, where $\alpha$ is the learning rate and $\lambda$ is the weight decay coefficient. The total KL loss can be approximated as:
$\mathcal{L}_{\text{kl}}\approx\|\phi^{t}-\phi^{t+1}\|^{2}=\alpha^{2}\|\nabla_{\phi}\mathcal{L}_{\text{original}}+\lambda\phi_{t}\|^{2}.$ (17)
When the training process converges to a minimum, the task-related gradients ($\nabla_{\phi}\mathcal{L}_{\text{original}}$) become small and nearly zero, and the KL loss becomes dominated by the weight decay term, and the approximation simplifies to: $\mathcal{L}_{\text{kl}}\approx\alpha^{2}\lambda^{2}\|\phi^{t}\|^{2}$. Thus, the L2 norm contributes to minimizing the KL loss. ∎
Next, we provide proof of the theorem for the M-Step.
Theorem A.2.
Assuming $q^{t+1}_{\phi_{t}}(Z)=p^{t+1}(Z)$ and an L2 norm is applied on $\phi$, the variational lower bound in Eq. (3) can be approximated as follows:
$\mathbb{E}_{p(X)}\left[\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\left[\log p^{t+1}_{\Theta_{t+1}}(\hat{Y}|X,Z)\right]\right]-C,$ (18)
which is equivalent to minimizing the expected reconstruction loss $\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{recon}(\mathbf{x},\hat{\mathbf{x}})]$
Proof.
In the M-step, $q_{\phi}(Z|X)$ serves as the estimator for the environment distribution $p^{t+1}(Z)$ defined by $q^{t+1}_{\phi_{t}}(Z)$ in the previous E-step. Then, we derive the expected log-likelihood for $\log p_{\Theta}(\hat{Y}|do(X))$ from Eq. (3) :
$\quad E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}$
$=\sum_{\mathbf{x}}\log{p_{\Theta}(\hat{Y}|do(X)=\mathbf{x})}p(X=\mathbf{x})$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}p(Z=z)$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}p(Z=z)\cdot\frac{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$\geq\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
(Jensen’s Inequality)
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{p_{\Theta}(\hat{Y}|X=\mathbf{x},Z=z)}$
$\hskip 58.32823pt+\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{\mathbf{z}}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
$=\underbrace{\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}}_{-\mathcal{L}_{recon}}$
$+\underbrace{\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}}_{-\mathcal{L}_{reg}}.$
The first term, similar to the justification in the E-step, maximizes the predictive power of the model in a batch-wise manner, while the second term serves as a regularization of the environment estimator. Since we assume $q^{t+1}_{\phi_{t}}(Z)=p(Z)$, we prove that the second term can be further reduced:
$-\mathcal{L}_{reg}$ (19)
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{\frac{p(Z=z)}{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}}$
$=\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\log{p(Z=z)}$
$\hskip 28.45274pt-\sum_{\mathbf{x}}p(X=\mathbf{x})\sum_{z}q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})\cdot\log{q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})}$
$=\sum_{z}\log{p(Z=z)}\sum_{\mathbf{x}}p(X=\mathbf{x})q^{t+1}_{\phi_{t}}(Z=z|X=\mathbf{x})+H_{\phi_{t}}(Z|X)$
$=\sum_{z}\log{q^{t+1}_{\phi_{t}}(Z=z)}q^{t+1}_{\phi_{t}}(Z=z)+H_{\phi_{t}}(Z|X)$
$=-H_{\phi_{t}}(Z)+H_{\phi_{t}}(Z|X)$
$=-I_{\phi_{t}}(Z;X),$
where $H_{\phi_{t}}(Z)$ and $H_{\phi_{t}}(Z|X)$ are the entropy and the conditional entropy induced by the estimator $q^{t+1}_{\phi_{t}}$ respectively, and $I_{\phi_{t}}$ denotes the mutual information. Since $I_{\phi^{t+1}}(Z;X)$ is approximated to a constant $C$ as $p_{s}(X)\approx p(X)$, we are able to acquire the final lower bound as:
$E_{p(X)}\log{p_{\Theta}(\hat{Y}|do(X))}\geq\mathbb{E}_{p(X)}\mathbb{E}_{q^{t+1}_{\phi}(Z|X)}\log{p_{\Theta}(\hat{Y}|X,Z)}-C$ (20)
Therefore we only focus on maximizing the first term, which becomes minimizing the reconstruction loss $\mathcal{L}_{recon}$. In practice, we make $q^{t+1}_{\phi}$ active during this step while posing an L2 norm on its parameters to better optimize the environment estimator as well as switching to a different environment distribution space. During training, as proven in Proof A.3.3, the updated $q^{t+1}_{\phi_{t+1}}$ eventually approximates the $q^{t+1}_{\phi_{t}}$.
A.4 Pre-train and Downstream Datasets Details
Due to incomplete records and limited non-U.S. data, we curated diverse, high-quality datasets emphasizing:
  1. Epidemic variation: Decades of weekly data (visualization) reflect multi-wave trends and interventions.
  2. Geographic spread: Covers 50 U.S. states and 25 non-U.S. Dengue regions.
  3. Climate diversity: Implicit variation (e.g., tropical Florida vs. temperate Maine).
  4. Data quality: Manually curated, gap-free datasets (from CDC, ProjectTycho, etc.).
We apply z-score normalization for all datasets to align disease data with zero-mean.
A.4.1 Pre-train Datasets
In this study, we utilize a comprehensive collection of 17 distinct diseases from the United States, sourced from Project Tycho. These diseases encompass both respiratory and non-respiratory categories and serve as the foundation for pre-training three transformer-based models: CAPE, PEM, and PatchTST. The selection criteria for these datasets were meticulously chosen based on the following factors:
Temporal Coverage and Geographic Representation: We prioritized diseases with extensive time series data and coverage across multiple regions to ensure the models are trained on diverse and representative datasets.
Consistent Sampling Rate: All selected datasets maintain a uniform sampling rate, which is crucial for the effective training of transformer models that rely on temporal patterns.
Data Quantity: Diseases with larger datasets in terms of both temporal length and the number of regions were preferred to enhance the robustness and generalizability of the models.
Among the 17 diseases, five are classified as non-respiratory, providing a balanced representation that allows the models to learn from varied disease dynamics. Before the pre-training phase, each disease dataset underwent a normalization process to standardize the data scales, ensuring comparability across different diseases. Subsequently, the datasets were aggregated at the national level based on their corresponding timestamps. The details of the pre-training datasets are summarized in Table 6.
Table 6: Pre-training Datasets collected from Project Tycho.
Disease Number of States Total Length Non-Respiratory
Gonorrhea 39 37,824 Yes
Meningococcal Meningitis 37 44,890 No
Varicella 30 33,298 No
Typhoid Fever 44 89,868 Yes
Acute Poliomyelitis 47 74,070 Yes
Hepatitis B 31 34,322 Yes
Pneumonia 41 68,408 No
Hepatitis A 38 37,303 Yes
Influenza 42 61,622 No
Scarlet Fever 48 129,460 No
Smallpox 44 71,790 No
Tuberculosis 39 95,564 No
Measles 50 151,867 No
Diphtheria 46 112,037 No
Mumps 41 50,215 No
Pertussis 46 109,761 No
Rubella 7 6,274 No
A.4.2 Downstream Datasets
In addition, we collect seven datasets of different types of diseases from diverse sources for downstream evaluations, which are all normalized without further processing. A summary of the downstream datasets is shown in Table 7.
Table 7: Statistics of the downstream datasets for evaluation.
Disease Number of Regions Sampling Rate Respiratory Total Length
ILI USA 1 Weekly Yes 966
ILI Japan 1 Weekly Yes 348
Measles 1 Bi-weekly Yes 1,108
Dengue 23 Mixed No 10,739
RSV 13 Weekly Yes 4,316
MPox 1 Daily No 876
COVID 16 Daily Yes 12,800
All collected diseases can be categorized into Respiratory and Non-respiratory types, which differ in their modes of transmission:
Respiratory. Respiratory diseases are transmitted primarily through the air via aerosols or respiratory droplets expelled when an infected individual coughs, sneezes, or talks. These diseases predominantly affect the respiratory system, including the lungs and throat.
Non-respiratory. Non-respiratory diseases are transmitted through various other routes such as direct contact, vectors (e.g., mosquitoes, ticks), contaminated food or water, and sexual activities. These diseases can affect multiple body systems and have diverse transmission pathways unrelated to the respiratory system.
A more detailed description of each dataset is shown below:
  1. ILI USA  (Centers for Disease Control and Prevention, 2023a): The weekly influenza-like-illness infection was reported by the CDC in the United States. We use the national-level infection counts from 2002 to 2020, which include various disease such as H1N1, H3N2v, etc.
  2. ILI Japan  (National Institute of Infectious Diseases, 2023): This dataset is collected from the Infectious Diseases Weekly Report (IDWR) in Japan, which contains national counts of weekly influenza-like-illness infections from August 2012 to March 2019.
  3. Measles  (Lau et al., 2020a): The measles dataset contains biweekly measles infections in England from 1906 to 1948.
  4. Dengue  (OpenDengue, 2023): OpenDenue aims to build and maintain a database of dengue case counts for every dengue-affected country worldwide since 1990 or earlier. We selected 23 countries for the experiment, which reports daily to weekly infections.
  5. RSV  (Centers for Disease Control and Prevention, 2023c): The Respiratory Syncytial Virus (RSV) infections in the US are reported by the RSV-NET from CDC. We use the weekly infections across 13 states from 2016 to 2024.
  6. MPox  (Centers for Disease Control and Prevention, 2023b): The clade II MPox case trends data in the US is reported by CDC. We use the nationwide weekly infections from 2022 to 2024.
  7. COVID  (Dong et al., 2020): The original data is from the Novel Coronavirus Visual Dashboard operated by the Johns Hopkins University Center for Systems Science and Engineering (JHU CSSE). We use the daily COVID-19 infections collected by JHU from 2020 to 2022 across 16 states.
A.5 Implementation Details
Motivation for Our Settings. Our primary focus is on pre-training epidemic forecasting models using temporal (time series) data rather than spatiotemporal data. This design choice is motivated by the following considerations:
  1. We aim to establish the foundation for epidemic pre-training in the temporal setting, which remains unexplored — only one prior work  (Kamarthi & Prakash, 2023) addresses this area, and it also focuses on temporal setting. As shown in Sections 4.2-4.5, we address critical questions around generalization, few-shot/zero-shot performance, and pre-training dynamics — topics that remain open even without spatial context.
  2. Temporal models are broadly applicable and more data-efficient, especially when spatial data is unavailable or unreliable. Many real-world epidemic datasets lack well-defined spatial graphs, and building them (e.g., from mobility or administrative data) is costly and complex, particularly at scale. These inconsistencies also hinder fair comparisons between temporal and spatiotemporal models.
  3. Our framework is extensible to spatiotemporal modeling. Specifically, the temporal input can be replaced with graph structured data, and the predictor can incorporate graph-based encoders. Exploring this direction is exciting future work, but we believe temporal pre-training is a crucial first step toward that goal.
Zero-Shot. Once pre-trained, our CAPE framework can be directly utilized for zero-shot forecasting where the model remains frozen and no parameter is updated. Similar to the MOMENT model (Goswami et al., 2024), we retain the pre-trained reconstruction head and mask the last patch of the input to perform forecasting: $\hat{\mathbf{y}}=\hat{\mathbf{x}}_{[T-c:T]}$.
Data Splits. For the ILI USA, Measles, and Dengue datasets, we split the data into 60% training, 10% validation, and 30% test. Other datasets are divided into 40% training, 20% validation, and 40% test. During test, we use the model checkpoint with the best validation performance.
Model Details. We design our model by stacking 4 layers of the CAPE encoder, each with a hidden size of 512 and 4 attention heads. For environment representations, we incorporate 16 distinct environments, each encoded with a size of 512. To ensure a fair comparison, PatchTST is configured with the same number of layers and hidden size as our CAPE-based model. For all other baseline models, we adopt the architectures as reported in previous studies (Wang et al., 2024a; Kamarthi & Prakash, 2023; Panagopoulos et al., 2021).
Training Details:
  1. We adopt an input length of 36 (Wu et al., 2023; Wang et al., 2024b) and a patch size of 4 for applicable models. For the environment estimator defined in Eq. (4), a shared weight $w_{k}$ is used for all environment representations. All results are evaluated using Mean Squared Error (MSE).
  2. CAPE follows the general EM framework, whose convergence is well-studied (Obata et al., 2024; Qu et al., 2019; Wu, 1983). To ensure stable convergence, we use a small learning rate (1e-5) with L2 regularization, train for 150 epochs, and select the model with the lowest validation error.
  3. For the training process, we pre-train CAPE, PEM, and PatchTST on a single Nvidia A100 GPU. During pre-training, we utilize only 70% of the available training data, specifically the first 70% of the dataset for each disease category. We set the learning rate to $1\times 10^{-5}$ and apply a dropout rate of 0.1 to prevent overfitting. In the CAPE pre-training strategy, we assign a weight of 0.5 to $\alpha$ to balance the contributions of contrast loss to the whole loss function. A detailed illustration of our pre-training strategy is shown in Algorithm 1.
  4. After pre-training, we fine-tune the entire model using a single Nvidia K80 GPU, maintaining the same hyperparameter settings for consistency. The best-performing model is selected based on its performance on the validation set. Similarly, for all baseline models, we train each until convergence and select the optimal model based on validation set performance for the subsequent test.
Fine-Tuning. For downstream tasks, we fine-tune the entire model using MSE loss, still employing EM for optimization. The difference is that we replace the task-specific head $h_{\psi}$ for pre-training, which is a linear transformation on each patch, with a head for forecasting that takes the concatenated latent representations $\mathbf{x}^{(L)}$ and maps them to the future prediction: $\hat{\mathbf{y}}=h_{\psi}(\mathbf{x}^{(L)})=\mathbf{W}\mathbf{x}^{(L)}$.
Algorithm 1 Hierarchical Contrastive Loss Optimization with EM
1:Input: $\mathcal{D}_{tr}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, $\alpha$, $\eta_{\theta,\psi,\phi}$, $\eta_{\mathbf{E}}$, $T_{max}$
2:Initialize: $\theta$, $\psi$, $\phi$, $\mathbf{E}\sim\text{Orthogonal}(\mathbf{I})$
3:repeat
4:    E-Step: Optimize Environment Representations $\mathbf{E}$
5:     Freeze $(\theta,\psi)$ and set $\alpha\leftarrow 0$
6:     Sample mini-batch $\mathcal{B}\subseteq\mathcal{D}_{tr}$
7:     Compute reconstruction loss: $\mathcal{L}_{recon}$
8:     Compute gradient: $\nabla_{\mathbf{E}}\mathcal{L}_{recon}$
9:     Update $\mathbf{E}$: $\mathbf{E}\leftarrow\mathbf{E}-\eta_{\mathbf{E}}\nabla_{\mathbf{E}}\mathcal{L}_{recon}$
10:    M-Step: Optimize Model Parameters $(\theta,\psi,\phi)$
11:     Freeze $\mathbf{E}$, set $\alpha$ to predefined values
12:     Sample mini-batch $\mathcal{B}\subseteq\mathcal{D}_{tr}$
13:     Compute contrastive loss: $\mathcal{L}_{CL}$
14:     Compute total loss: $\mathcal{L}_{final}=\mathcal{L}_{recon}+\alpha\mathcal{L}_{CL}$
15:     Update parameters:
16:       $(\theta,\psi,\phi)\leftarrow(\theta,\psi,\phi)-\eta_{\theta,\psi,\phi}\nabla_{\theta,\psi,\phi}\mathcal{L}_{final}$
17:until $t=T_{max}$
18:Output: $(\theta,\psi,\phi),\mathbf{E}$
A.6 Evaluation Metrics
In this study, we primarily use MSE for its sensitivity to hard cases (large errors) and common use in epidemic forecasting  (Datilo et al., 2019). In addition, we also report MAPE and peak timing error (averaged over horizons 1–16) on the ILI USA dataset as an example:
Table 8: Model Performance Metrics
Model MAPE Peak Error
CAPE 0.188 1.226
Informer 0.246 1.553
LSTM 0.264 1.398
DLinear 0.225 1.442
A.7 Environment Influence
We also add Gaussian noise to environment weights (Equation 4), causing performance drops, which highlights the importance of accurate environment estimation:
Table 9: Model Performance under Different Noise Scales
Noise Scale (%) 5% 10% 15% 20%
MSE 0.327 0.373 0.456 0.559
MAPE 0.198 0.211 0.236 0.264
A.8 Full Results on Pre-train datasets
In addition to evaluating the performance of the models on downstream datasets, we also provide the in-domain evaluation results from the pre-training datasets. Recall that we used 70% data of each disease for pre-training, here we fine-tuned the model on the 70% of each disease and evaluate both CAPE and the pre-trained PatchTST on the rest 30% data. As shown in Table 10, CAPE consistently outperforms PatchTST on 13/15 datasets, proving the effectiveness of our method.
Table 10: Performance of CAPE and pre-trained PatchTST across diseases in the pre-training datasets. The results presented is the average over horizons of 1,2,4,8,16.
Disease Method Horizon 1 Horizon 2 Horizon 4 Horizon 8 Horizon 16 Average
Mumps CAPE 0.000284 0.000290 0.000370 0.000451 0.000539 0.000387
PatchTST 0.000280 0.000310 0.000388 0.000508 0.000627 0.000423
Meningococcal Meningitis CAPE 0.063022 0.066196 0.073552 0.093547 0.108842 0.081032
PatchTST 0.054611 0.061641 0.073794 0.088404 0.096449 0.074980
Influenza CAPE 0.367677 0.510453 0.693110 0.903920 1.037177 0.702467
PatchTST 0.392925 0.644013 0.717147 0.851498 1.061066 0.733330
Hepatitis B CAPE 0.071834 0.072827 0.074606 0.077816 0.068012 0.073019
PatchTST 0.074016 0.082576 0.084535 0.085867 0.074103 0.080219
Pneumonia CAPE 0.038916 0.052092 0.082579 0.137004 0.191675 0.100453
PatchTST 0.036961 0.074596 0.096963 0.152206 0.174871 0.107119
Typhoid Fever CAPE 0.004918 0.004393 0.004552 0.005051 0.005828 0.004948
PatchTST 0.007068 0.005954 0.005906 0.006519 0.006709 0.006431
Hepatitis A CAPE 0.347792 0.349403 0.352361 0.360705 0.315496 0.345151
PatchTST 0.331339 0.349549 0.356113 0.381637 0.338067 0.351341
SCAPEet Fever CAPE 4.229920 5.258288 6.787577 10.865951 13.724634 8.173274
PatchTST 8.561295 13.564009 17.241462 19.315905 20.373520 15.811238
Gonorrhea CAPE 0.010826 0.010900 0.011246 0.011483 0.011898 0.011271
PatchTST 0.011297 0.012223 0.013411 0.013438 0.013241 0.012722
Smallpox CAPE 0.063829 0.065191 0.076199 0.098973 0.157850 0.092408
PatchTST 0.070972 0.076843 0.107076 0.124042 0.165442 0.108875
Acute Poliomyelitis CAPE 0.254014 0.394454 0.355898 0.480525 0.745428 0.446064
PatchTST 0.094695 0.134304 0.270908 0.392511 0.482426 0.274969
Diphtheria CAPE 0.006789 0.005360 0.006557 0.010682 0.014136 0.008705
PatchTST 0.011019 0.008891 0.009036 0.013048 0.015531 0.011505
Varicella CAPE 0.000119 0.000128 0.000154 0.000212 0.000245 0.000171
PatchTST 0.000109 0.000141 0.000169 0.000237 0.000296 0.000190
Tuberculosis CAPE 0.178741 0.170441 0.215367 0.177671 0.198068 0.188057
PatchTST 0.189156 0.209008 0.189944 0.204680 0.277632 0.214084
Measle CAPE 0.009626 0.010982 0.016451 0.022407 0.042980 0.020489
PatchTST 0.013008 0.012608 0.020903 0.039835 0.063844 0.030039
A.9 Full results for few-shot forecasting
We present the complete few-shot performance across different horizons in Table 11. While CAPE does not achieve state-of-the-art average performance on the ILI USA dataset with limited training data, it excels in short-term forecasting when the horizon is smaller.
Table 11: Few-shot learning results with horizons ranging from 1 to 16 future steps. The length of the lookback window is set to 36. Each model is evaluated after being trained on 20%, 40%, 60% and 80% of the full training data.
Dataset Horizon CAPE PatchTST Dlinear MOMENT PEM
20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100% 20% 40% 60% 80% 100%
ILI USA 1 1.155 0.535 0.307 0.178 0.155 1.361 0.662 0.355 0.191 0.195 1.430 1.000 0.460 0.230 0.170 2.859 1.274 0.608 0.267 0.216 1.424 0.620 0.330 0.189 0.145
2 1.396 0.925 0.465 0.220 0.200 1.389 0.806 0.489 0.234 0.264 2.210 1.090 0.660 0.280 0.220 3.242 1.709 0.695 0.342 0.271 1.463 0.829 0.434 0.256 0.210
4 1.770 1.154 0.640 0.306 0.270 1.923 1.215 0.656 0.387 0.385 2.500 1.670 0.720 0.380 0.310 3.910 1.901 0.891 0.399 0.356 1.889 1.186 0.625 0.393 0.312
8 2.611 1.912 0.978 0.519 0.404 2.713 1.623 0.833 0.544 0.535 3.510 1.970 0.980 0.530 0.450 4.706 2.013 1.120 0.615 0.482 2.649 1.690 0.966 0.580 0.573
16 3.674 2.473 1.411 0.622 0.516 3.182 1.789 1.056 0.649 0.485 4.460 2.240 1.260 0.640 0.580 5.233 2.335 1.251 0.669 0.580 3.294 1.979 1.049 0.679 0.526
Avg 2.121 1.400 0.760 0.369 0.309 2.114 1.219 0.677 0.401 0.373 2.822 1.594 0.816 0.412 0.346 3.990 1.847 0.913 0.459 0.381 2.143 1.261 0.681 0.419 0.353
Dengue 1 3.254 1.384 0.489 0.384 0.218 3.700 1.580 0.657 0.389 0.203 3.600 1.470 0.550 0.350 0.220 4.585 2.480 0.689 0.423 0.383 3.383 1.613 0.558 0.350 0.206
2 4.463 2.340 0.735 0.487 0.301 5.832 2.159 0.846 0.507 0.296 7.090 2.170 0.820 0.510 0.310 6.609 2.990 0.922 0.587 0.521 5.404 2.257 0.869 0.507 0.300
4 7.563 3.728 1.250 0.817 0.540 9.525 3.636 1.517 1.069 0.588 11.190 4.130 1.520 0.940 0.560 12.877 4.106 1.644 0.966 0.669 8.782 4.428 1.608 1.037 0.522
8 15.526 7.276 2.836 1.922 1.193 19.052 9.530 3.597 2.133 1.296 21.910 9.690 3.470 2.160 1.250 23.298 9.229 3.625 2.135 1.235 17.023 8.117 3.323 2.249 1.295
16 35.870 17.204 6.469 3.946 2.210 30.451 19.616 7.238 4.289 2.536 35.350 24.640 7.890 4.780 3.060 31.115 18.877 7.200 4.551 3.984 29.934 18.861 7.368 4.390 2.497
Avg 13.335 6.386 2.356 1.511 0.892 13.712 7.304 2.771 1.678 0.984 15.828 8.420 2.850 1.748 1.080 15.697 7.536 2.816 1.733 1.358 12.90 7.055 2.745 1.707 0.964
Measles 1 0.168 0.158 0.107 0.095 0.069 0.400 0.217 0.121 0.091 0.094 0.560 0.470 0.190 0.150 0.100 1.211 0.316 0.138 0.108 0.102 0.227 0.200 0.106 0.106 0.084
2 0.229 0.256 0.165 0.134 0.096 0.511 0.325 0.186 0.148 0.127 0.680 0.400 0.320 0.220 0.150 1.376 0.367 0.159 0.167 0.138 0.313 0.339 0.155 0.153 0.127
4 0.371 0.399 0.267 0.198 0.155 0.663 0.510 0.297 0.243 0.205 1.050 0.920 0.360 0.310 0.240 1.444 0.516 0.278 0.228 0.196 0.497 0.451 0.258 0.240 0.196
8 0.564 0.776 0.451 0.339 0.280 1.050 1.269 0.479 0.414 0.378 1.580 1.340 0.660 0.540 0.450 1.895 1.181 0.507 0.386 0.883 0.865 1.213 0.487 0.441 0.382
16 1.086 1.408 0.917 0.658 0.743 1.692 1.847 1.157 0.900 0.723 2.100 2.520 1.480 1.170 1.030 2.379 2.192 1.041 1.468 1.183 1.448 2.275 1.145 0.880 0.740
Avg 0.483 0.600 0.381 0.285 0.269 0.863 0.834 0.448 0.359 0.306 1.194 1.130 0.602 0.478 0.394 1.661 0.915 0.425 0.471 0.500 0.670 0.896 0.430 0.364 0.306
A.10 Impact of the Environment Distribution and Pre-train Epochs
Figure 9: Downstream performance vs. environment distribution and pre-train epochs.
A.11 Impact of pre-train ratio on the downstream datasets
We provide additional evaluations for CAPE on downstream datasets to analyze the impact of the pre-training ratio. As shown in Figure LABEL:fig:pretrain_scale, increasing the pre-training ratio eventually improves downstream performance across all datasets.
(a) Dengue
(b) Measles
(c) ILI USA
(d) COVID
Figure 10: Downstream performance with different ratios of pre-training datasets. The input length is set to 36 and all MSE results are averaged over {1,2,4,8,16} future steps.
A.12 Mitigating Distribution Shift
Figure 11: CMD scores w/ and w/o environment estimate between train/test sets. We measure the CMD scores based on the last embedding output by our model.
A.13 Disentangling Disease-Specific Dynamics
After pre-training, both CAPE and PatchTST were frozen and applied to downstream datasets directly to generate latent space embeddings for each sample. We then employed the Davies-Bouldin Index (DBI) to assess the separability of these embeddings by disease. CAPE consistently achieved lower DBI scores (we show an example in Figure 16), demonstrating a superior ability to distinguish between different diseases in the latent space compared to PatchTST. This enhanced separability highlights the effectiveness of CAPE’s environment estimation and backdoor adjustment. These mechanisms are crucial for mitigating the confounding influence of noisy, spurious, and environment-dependent factors ($X_{sp}$), thereby enabling the model to better isolate and represent the unique, intrinsic, and causal disease dynamics ($X_{ca}$).
Figure 12: DBI between the embeddings of each pair of downstream datasets from the pre-trained model. (See Figure  16).
A.14 Visualization of the Estimated Environments
According to $\hat{\mathbf{e}}^{(l)}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k}^{(l)}$, an aggregated environment is the weighted sum of the learned latent environment representations. Therefore, the estimation shares the same latent space as the fixed representations and we are able to visualize them using t-SNE. As shown in Figure 13, we visualize the aggregated environments (Estimated) as well as the learned latent environment (Anchor) from a CAPE model with 8 environments.
(a) ILI USA
(b) Measles
(c) Dengue
Figure 13: Visualization of the estimated environment representations using t-SNE.
A.15 Visualization of Distribution Shift for Downstream Datasets
We provide a visualization of the sample distribution used in this study. Each sample has a fixed length of 36, representing the historical infection trajectory. To better understand the distributional differences, we use t-SNE to reduce the data to one dimension and visualize the training and test samples using different colors. As shown in Figure 14, a significant distribution shift is visually apparent across most datasets. To quantitatively assess the distributional differences between the training and test sets, we calculate the Central Moment Discrepancy (CMD) score (Zellinger et al., 2017). The CMD score measures the discrepancy between the central moments of the two distributions up to a specified order $K$. For two distributions $X$ (training set) and $X_{\text{test}}$ (test set), the CMD score is defined as:
$\quad\text{CMD}(X,X_{\text{test}})=\|\mu_{1}(X)-\mu_{1}(X_{\text{test}})\|_{2}+\sum_{k=2}^{K}\|\mu_{k}(X)-\mu_{k}(X_{\text{test}})\|_{2},$ (21)
where:$\mu_{k}(X)$ denotes the $k$-th central moment of $X$, defined as: $\mu_{k}(X)=\mathbb{E}[(X-\mathbb{E}[X])^{k}],$ and similarly for $\mu_{k}(X_{\text{test}})$. $\|\cdot\|_{2}$ is the Euclidean norm. $K$ is the maximum order of moments considered. The CMD score aggregates the differences in the mean (first moment) and higher-order moments (e.g., variance, skewness), providing a robust measure of the distribution shift. In our experiments, we set $K=3$ to capture up to the third-order central moments. This score quantitatively complements the visual observations in Figure 14, offering a more comprehensive understanding of the distributional differences between training and test sets.
Impact of Distribution Shifts. Distribution shifts between training and test datasets pose significant challenges to the generalizability and robustness of predictive models. When the underlying data distributions differ, models trained on the training set may struggle to maintain their performance on the test set, leading to reduced accuracy and reliability. These discrepancies can arise from various factors, such as temporal changes in infection patterns or geographical variations. In this paper, we assume that the inherent infection pattern of a particular disease remains constant, and the distribution shifts for the disease are primarily caused by the rapidly changing environment, which results in diverse infection patterns. In the context of epidemic modeling, such shifts are especially critical, as they can undermine the model’s ability to accurately predict future infection trends, which is essential for effective public health interventions.
Figure 14: The KDE plot of training set and test set. Each sample contains an infection trajectory of 36 weeks. t-SNE is applied to visualize the distributions of both sets.
A.16 Latent space visualization of Measle and Covid datasets from pre-trained models.
In order to demonstrate that CAPE effectively disentangles the underlying dynamics of diseases from the influence of the environment, we visualize the output embeddings for the Measles and COVID datasets by projecting them into a two-dimensional space using t-SNE. Specifically, we utilize the pre-trained model without fine-tuning on these two downstream datasets and visualize $\mathbf{x}^{(L)}$, the final-layer embeddings, as individual data points in the figure. As shown in Figure 16, CAPE (left) visually separates the two datasets more effectively than the pre-trained PatchTST model (right). To quantitatively evaluate the separability of the embeddings, we compute the Davies–Bouldin Index (DBI), which is defined as:
$\text{DBI}=\frac{1}{K}\sum_{i=1}^{K}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{\|\mu_{i}-\mu_{j}\|}\right),$ (22)
where $K$ is the number of clusters (in this case, two: Measles and COVID), $\mu_{i}$ is the centroid of cluster $i$, $\sigma_{i}$ is the average intra-cluster distance for cluster $i$, defined as: $\sigma_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}\in C_{i}}\|\mathbf{x}-\mu_{i}\|$, where $C_{i}$ is the set of points in cluster $i$, $\|\mu_{i}-\mu_{j}\|$ is the Euclidean distance between the centroids of clusters $i$ and $j$. The DBI measures the ratio of intra-cluster dispersion to inter-cluster separation. Lower DBI values indicate better separability. As shown in Figure 16, CAPE achieves a significantly lower DBI compared to PatchTST, confirming its superior ability to disentangle the underlying disease dynamics from environmental factors. A more complete result is shown in Figure 12.
A.17 Hyper-parameter Sensitivity Analysis
While we treat $\hat{\mathbf{E}}_{j,c}^{(l)}=\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}$ in the contrastive loss and aligned the combined representation of the input and environment, we further applied the contrastive loss on the aggregated environment representations at each layer, denoted as $\tilde{{\mathbf{E}}}_{j,c}^{(l)}=\hat{\mathbf{e}}^{(l)}_{c}$ and assigned a hyper-parameter $\beta$ to control for its weight:
$\mathcal{L}_{\text{final}}$ $=\sum_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}^{\prime}}^{(L)})+\frac{\beta}{L}\sum_{l}\mathcal{L}_{\text{CL}}(\tilde{{\mathbf{E}}}^{(l)},\tilde{{\mathbf{E}}^{\prime}}^{(l)}),$ (23)
$\hskip 85.35826pt\mathbf{X}\sim\mathcal{D}^{\prime}_{s}\in\mathcal{D}_{\text{pre}}$
where $L$ denotes the number of layers. Using Equation (23), we further fine-tune our pre-trained model by varying $\alpha$ and $\beta$ across the values [1e-3, 1e-2, 1e-1, 1, 10]. The results of this sensitivity analysis are presented in Figure 15.
(a) ILI USA
(b) Measles
(c) ILI Japan
Figure 15: Hyperparameters sensitivity of $\alpha$ and $\beta$.
Figure 16: Output latent space of two pre-trained models without fine-tuning from Measle and Covid datasets. Left: CAPE; Right: Pre-trained PatchTST.